# Sharp one-point estimates and Minkowski content for the scaling limit of three-dimensional loop-erased random walk

### v1

@@ 6.3 Invariance
For any $a>0$, let ${\cal K}^{a}$ be the scaling limit [^3] Its existence has also been confirmed in [10]. of 3D LERW of mesh size $m^{-1}$, $m\in{\mathbb{Z}^{+}}$ in $a\mathbb{D}$. Write $\eta^{a}$ for the parametrization of ${\cal K}^{a}$ by its Minkowski content. Then it follows that
@@
As it was discussed above in (2.19), Subsection 6.1 in [10] shows that ${\cal K}$ is both scale-invariant and rotation-invariant. Then ${\cal K}^{a}\overset{\rm d}{=}a{\cal K}$ and $T_{\theta}({\cal K})\overset{\rm d}{=}{\cal K}$. The the corresponding invariance of the Minkowski content and the definition of the parametrization by the Minkowski content in (1.6) imply the conclusion of the corollary. ∎

### v2

@@ 6.3 Invariance
For any $a>0$, let ${\cal K}^{a}$ be the scaling limit [^3] Its existence has also been confirmed in [11]. of 3D LERW of mesh size $^{-1}, m+m\in{\mathbb{Z}^{+}}$ in $a\mathbb{D}$. Write $\eta^{a}$ for the parametrization of ${\cal K}^{a}$ by its Minkowski content. Then it follows that
@@
As discussed above in (2.19), Subsection 6.1 in [11] shows that $\cal K$} is both scale-invariant and rotation-invariant. Then ${\cal K}^{a}\overset{\rm d}{=}a{\cal K}$ and $T_{\theta}({\cal K})\overset{\rm d}{=}{\cal K}$. The the corresponding invariance of the Minkowski content and the definition of the parametrization by the Minkowski content in (1.6) imply the conclusion of the corollary. ∎
Acknowledgments
The authors would like to thank the anonymous referees for very detailed and helpful comments.
SHT would like to acknowledge the support of UNAM-PAPIIT grant IA103724. XL is supported by National Key R&D Program of China (No. 2021YFA1002700 and No. 2020YFA0712900) and NSFC (No. 12071012). DS was supported by JSPS Grant-in-Aid for Scientific Research (C) 22K03336, JSPS Grant-in-Aid for Scientific Research (B) 22H01128 and 21H00989.